3. An object's velocity as a function of time is s
Answer #1

3)

a) total displacement = Area under v-t curve

= 1*2 + (1/2)*2*2 +2*3 + (1/2)*2*3 - (1/2)*1*3

= 11.5 m

b) position = 10 + 11.5

= 21.5 m

c)

total distance travelled = 1*2 + (1/2)*2*2 +2*3 + (1/2)*2*3 + (1/2)*1*3

= 14.5 m

d)

at t = 1s,

a = (v2-v1)/(t2-t1) = (3-1)/(2-1) = 2 m/s^2

at t = 3s,

a = (v2-v1)/(t2-t1) = (3-3)/(4-2) = 0 m/s^2

at t = 5s,

a = (v2-v1)/(t2-t1) = (0-3)/(6-4) = -1.5 m/s^2

at t = 6.5 s

a = (v2-v1)/(t2-t1) = (-3 - 0))/(7-6) = -3 m/s^2